# DeepResearchGym: A Free, Transparent, and Reproducible Evaluation Sandbox for Deep Research

### v2

@@ 3.2 Deep Research Evaluation Methods
To demonstrate how DeepResearchGym can support evaluation of deep research systems, we instantiate an evaluation protocol built around the Researchy Questions dataset [28]. This dataset was derived from Bing search logs, and contains complex queries associated with user engagement, making it well-suited for benchmarking systems that synthesize long-form, multi-perspective reports. While the sandbox is agnostic to the specific evaluation task and compatible with a broad range of use cases, we introduce this protocol to fill a current gap in the evaluation landscape, and to provide initial empirical observations using our API. Our protocol scores generated the reports along three axes: overall report quality, factual faithfulness to retrieved sources, and alignment with user information needs.
@@ 3.2.1 Researchy Questions
Evaluating deep research systems requires queries that naturally drive extensive information exploration and synthesis. The Researchy Questions dataset [28] was curated specifically to capture such queries. Rather than featuring simple factoid questions, the dataset consists of approximately 96,000 real-world information-seeking queries that led users to engage with multiple documents during search sessions, as measured by aggregated click distributions over ClueWeb22. For reference, Appendix A shows a sample of queries together with clicked document URLs.
The heavy engagement with diverse sources reflects the essential challenges deep research systems are designed to address: synthesizing information across multiple perspectives, reconciling conflicting evidence, and constructing comprehensive responses. In contrast to short, answerable queries that focus on locating a single fact, Researchy Questions exhibit the complexity that require broader exploration and deeper understanding. While the dataset was originally introduced for evaluating retrieval performance, its properties make it a strong foundation for studying long-form generation grounded in multi-document evidence. In the next section, we describe our evaluation methodology for extending the use of Researchy Questions to benchmark deep research generation.
@@ 3.2.2 Long-form Report Evaluation Metrics
Deep research systems focused on providing report-like answers face multiple challenges inherent to long-form generation evaluation [41], where outputs must be assessed not only for linguistic fluency and informativeness, but also for factual grounding and content relevance. We follow a tri-faceted evaluation framework that assesses the alignment with user information needs, factual grounding, and overall quality of generated answers. The Appendices contains all the prompts used for LLM-based metrics (Appendix B), an example report (Appendix C), and its detailed evaluation (Appendix D).
Report Relevance: As the primary metric for assessing user satisfaction, we evaluate how well the generated reports address the user’s underlying information needs. Given that Researchy Questions are derived from real-world web search sessions, we leverage the set of documents clicked by users as a proxy for ground-truth information targets. Following the Key Point Recall (KPR) methodology [26], we extract salient points from each ground-truth document using an LLM guided by structured prompts, capturing the core content users engaged with. We then assess each generated report for semantic inclusion of these key points, computing the KPR score as $\frac{1}{M}\sum_{j=1}^{M}c_{j}$, where $M$ is the total number of key points and $c_{j}$ indicates whether key point $j$ is supported by the report, as judged by an LLM.
@@
Retrieval Faithfulness: Beyond relevance, we assess the factual grounding of generated reports, adapting the LLM-as-a-judge approach of the TREC-RAG evaluation process [37]. Our automatic citation evaluation pipeline follows a three-stage process. First, factual claims are extracted from the report, along with any URLs referenced as support. Second, the content of each cited source is retrieved. Third, an LLM is prompted to assess whether the cited source adequately supports the corresponding claim. This procedure captures both the presence of citations and their substantive validity.
@@
Report Quality: To capture aspects of writing quality and analytical depth, we employ another LLM-as-a-Judge protocol [6], prompting a strong LLM to evaluate each answer along two key dimensions: clarity, reflecting logical coherence and linguistic fluency; and insightfulness, capturing analytical nuance and the depth of reasoning presented. These dimensions are commonly used in long-form generation evaluation [15, 31] and provide complementary evidence of the overall presentation quality of the generated content.
@@ 4 Benchmarking Deep Research Systems
Table 2: Comparison of deep research systems on the Researchy Questions test set using (i) each system’s original commercial search API and (ii) DeepResearchGym’s search API. Scores are judged by gpt-4.1-mini-2025-04-14. Systems marked with * are not tailored for long-report generation.
@@
System Commercial DRGym Commercial DRGym Commercial DRGym
@@
4.1 Experimental Setup
To evaluate the current landscape of deep research systems, we conducted a systematic benchmarking study, following the protocol described in Section 3.2.2 with gpt-4.1-mini-2025-04-14 as the LLM judge. We use a subset of the previously introduced Researchy Questions dataset, namely the top 1,000 queries from the test set, ranked by the number of documents clicked during the original search sessions. This ranking naturally favors queries that drive extensive exploration, aligning with the goals of deep research systems.
We evaluate a diverse set of deep research systems spanning both commercial and open-source implementations. The commercial systems include gpt4-search-preview from OpenAI and sonar-deepresearch from Perplexity, which represent the strongest variants available through the respective APIs. On the open-source side, we include GPT-Researcher and HuggingFace DeepSearch. All four systems are capable of generating long-form reports. We also evaluate three academic systems. OpenDeepSearch produces similarly comprehensive outputs, while Search-o1 and Search-R1 focus on concise, short-form answers. Although not designed for deep research tasks, these last two systems serve as lower-bound references and help verify that our evaluation metrics capture meaningful differences in generative capabilities. All systems are evaluated using their default configurations.
Experiments leveraging DeepResearchGym’s search API default to the ClueWeb22-B corpus, given the higher alignment with the Researchy Questions benchmark. Recall that, as discussed in Section 3.1.3, FineWeb is also available through our API endpoints, and, unlike ClueWeb22, this does not require signing the research-use license.
4.2 System-level Evaluation
@@
Third, significant performance gaps exist in citation practices. Systems like GPT-Researcher and OpenDeepSearch demonstrate strong citation precision and recall, effectively grounding factual claims in retrieved evidence. In contrast, systems such as HuggingFace’s deepresearch exhibit near-zero citation metrics, stemming from fundamental design choices that omit explicit citation mechanisms. This highlights a key distinction in how different approaches conceptualize evidence grounding, whether through explicit source attribution or implicit knowledge integration.
The evaluation also reveals an interesting trade-off in commercial systems, which tend to achieve excellent narrative quality but sometimes at the expense of granular citation precision. Manual inspection indicates two common patterns: (1) these systems often use citations to justify broader report sections rather than specific claims, and (2) they sometimes reference URLs that cannot be fully crawled. This suggests a potential tension between optimizing for narrative coherence versus precise evidence anchoring in current system designs.
@@ 4.3 Query-level Analysis
Figure 2: Query-level correlation between commercial search APIs and DeepResearchGym API evaluation metrics. The empty plot reflects a non-citation system.
To further investigate the consistency of system performance across individual queries, we conducted a fine-grained analysis comparing results obtained under each system’s original retrieval API, against those those from DeepResearchGym’s API, focusing only on the systems explicitly geared towards long-report generation. Figure 2 presents scatter plots of per-query scores across our three evaluation axes. The Spearman correlation coefficients quantify the alignment between retrieval configurations.
@@ 4.4 Human Evaluation
To validated our automatic evaluation protocol and assess its alignment with human judgments, we conduct a systematic human evaluation study. The evaluation set comprises 210 queries with their corresponding generated reports. For each query, we present annotators with two reports generated by different systems, asking them to select the better one based on overall informativeness, coherence, and factual accuracy. Seven annotators (all also co-authors of this paper) participated in the study, with each query independently evaluated by three different individuals. To mitigate potential bias, the evaluation was conducted double-blind: annotators were unaware of which systems produced each report, and system assignments were randomized across queries. Report presentation order was also randomized to prevent positional bias, and ties were disallowed to enforce binary preferences.
Table 3: LLM versus human agreement (Cohen’s $\kappa$).
Metric $\kappa$
KPR 0.72
Citation Precision 0.86
Citation Recall 0.86
Clarity 0.89
Insightfulness 0.84
Figure 3: System win rates as judged by human preferences.
To assess inter-annotator reliability, we compute pairwise Cohen’s $\kappa$ scores across all annotated comparisons. The average pairwise $\kappa$ is 0.87, indicating a high level of agreement and supporting the consistency of human judgments. Building on this, Table 4.4 presents Cohen’s $\kappa$ values measuring agreement between LLM-based automatic judgments and human preferences across each evaluation dimension, while Figure 3 presents pairwise win rates, showing the percentage of times each system was preferred over another based on human annotations. The KPC metric was excluded from the analysis in Table 4.4 due to a high rate of LLM-judged ties, leaving insufficient samples for analysis. The results demonstrate substantial agreement and the same trend regarding relative model performance. This validates our LLM-as-a-judge evaluation protocol, strengthening confidence in the benchmark results presented in the previous sections.
@@ 5 Conclusion and Future Work
DeepResearchGym offers a reproducible sandbox for developing and benchmarking deep research systems, providing a stable alternative to commercial search APIs. By anchoring retrieval to high-quality web corpora and combining the search API with a multi-dimensional evaluation protocol, our framework enables controlled benchmarking of systems that generate long-form reports.
@@
Our results demonstrate that DeepResearchGym’s search API provides a reliable research-grade substitute for commercial retrieval infrastructures. Across multiple evaluation axes, all evaluated systems maintain comparable performance when transitioning from proprietary search APIs to our transparent and reproducible retrieval environment, affirming that DeepResearchGym preserves the retrieval fidelity necessary to support complex report-style synthesis. By isolating system behavior from fluctuating retrieval conditions, DeepResearchGym offers a stable foundation for reproducible and fair benchmarking of deep research systems.
@@
Future extensions to DeepResearchGym can perhaps expand the coverage to larger and more recent web corpora, enabling evaluation of time-sensitive queries and emerging topics. Moreover, the integration of domain-specific benchmarks may further support assessment in high-stakes contexts such as healthcare or law, where retrieval precision and factual reliability are critical. Finally, the platform opens space for testing new agent architectures that incorporate structured planning, iterative search, and reinforcement learning, advancing research on deep research systems.
Limitations: Current limitations include: the reliance of deep research systems on proprietary and evolving LLMs for generation, which restricts full output reproducibility; evaluation variance introduced by LLM-as-a-judge protocols despite prompt/output standardization; and the use of static web corpora, which limits coverage of emerging or time-sensitive content. Ethical risks also apply, such as amplification of bias, misinformation, or misuse of retrieved content, although DeepResearchGym is targeted at academic research, instead of real-world deployments.
Acknowledgments and Disclosure of Funding
Appendix A Researchy Questions

### v3

@@ 3.2 Deep Research Evaluation Methods
To demonstrate how DeepResearchGym can support evaluation of deep research systems, we instantiate an evaluation protocol built around the Researchy Questions dataset (DBLP:journals/corr/abs-2402-17896). While the sandbox is agnostic to the specific evaluation task and compatible with a broad range of use cases, we introduce this protocol to fill a gap in the evaluation landscape, and to provide initial empirical observations using our API.
@@ 3.2.1 Researchy Questions
Evaluating deep research systems requires queries that naturally drive extensive information exploration and synthesis. The Researchy Questions dataset (DBLP:journals/corr/abs-2402-17896) was curated specifically to capture such queries. Rather than featuring simple factoid questions, the dataset consists of approximately 96,000 real-world information-seeking queries that led users to engage with multiple documents during search sessions, as measured by aggregated click distributions over ClueWeb22. For reference, Appendix B shows a sample of queries together with clicked document URLs, as well as an analysis of query time sensitivity.
The heavy engagement with diverse sources reflects the essential challenges deep research systems are designed to address: synthesizing information across multiple perspectives, reconciling conflicting evidence, and constructing comprehensive responses. While the dataset was originally introduced for evaluating retrieval performance, its properties make it a strong foundation for studying long-form generation grounded in multi-document evidence. In the next section, we describe the proposed evaluation methodology, which extends the use of Researchy Questions to benchmark deep research generation.
@@ 3.2.2 Long-form Report Evaluation Metrics
Deep research systems that focus on providing report-like answers face multiple challenges inherent to long-form generation evaluation (DBLP:conf/acl/XuSIC23), where outputs must be assessed not only for linguistic fluency and informativeness, but also for factual grounding and content relevance. We follow a tri-faceted evaluation framework that assesses the alignment with user information needs, factual grounding, and overall quality of generated answers. The Appendices contains all the prompts used for LLM-based metrics (Appendix C), an example report (Appendix D), and its detailed evaluation (Appendix E).
Report Relevance: As the primary metric for assessing user satisfaction, we evaluate how well the generated reports address the user’s underlying information needs. Given that Researchy Questions are derived from real-world web search sessions, we leverage the set of documents clicked by users as a proxy for ground-truth information targets. Following the Key Point Recall (KPR) methodology (DBLP:conf/emnlp/QiXGWZX24), we extract salient points from each ground-truth document using an LLM guided by structured prompts, capturing the core content users engaged with. We then assess each generated report for semantic inclusion of these key points, computing the KPR score as $\frac{1}{M}\sum_{j=1}^{M}c_{j}$, where $M$ is the total number of key points and $c_{j}$ indicates whether key point $j$ is supported by the report, as judged by an LLM.
@@
Retrieval Faithfulness: Beyond relevance, we assess the factual grounding of generated reports, adapting the LLM-as-a-judge approach of the TREC-RAG evaluation process (thakur2025supportevaluationtrec2024). Our automatic citation evaluation pipeline follows a three-stage process. First, factual claims are extracted from the report, along with any URLs referenced as support. Second, the content of each cited source is retrieved. Third, an LLM is prompted to assess whether the cited source adequately supports the corresponding claim. This procedure captures both the presence of citations and their substantive validity.
@@
Report Quality: To capture aspects of writing quality and analytical depth, we employ an LLM-as-a-Judge protocol (DBLP:journals/corr/abs-2411-15594), prompting a strong LLM to evaluate each answer along two key dimensions: clarity, reflecting logical coherence and linguistic fluency; and insightfulness, capturing analytical nuance and the depth of reasoning presented. These dimensions are commonly used in long-form generation evaluation (DBLP:conf/emnlp/LiuIXWXZ23; DBLP:conf/naacl/SahaLCBWL24) and provide evidence of the presentation quality of the generated content.
@@ 4 Benchmarking Deep Research Systems
4.1 Experimental Setup
To evaluate the current landscape of deep research systems, we conducted a systematic benchmarking study, following the protocol described in Section 3.2.2 with gpt-4.1-mini-2025-04-14 as the LLM judge. We used a subset of the previously introduced Researchy Questions dataset, namely the top 1,000 queries from the test set, ranked by the number of documents clicked during the original search sessions. This ranking naturally favors queries that drive extensive exploration, aligning with the goals of deep research systems.
We evaluated a diverse set of deep research systems spanning both commercial and open-source implementations. The commercial systems include gpt4-search-preview from OpenAI and sonar-deepresearch from Perplexity, which represent the strongest variants available through the respective APIs (at the time of writing). On the open-source side, we include GPT-Researcher and HuggingFace DeepSearch. All four systems are capable of generating long-form reports. We also evaluate three academic systems. OpenDeepSearch produces similarly comprehensive outputs, while Search-o1 and Search-R1 focus on concise, short-form answers. Although not designed for deep research tasks, these last two systems serve as lower-bound references and help verify that our evaluation metrics capture meaningful differences in generative capabilities.
All systems are evaluated in their default configurations, and DeepResearchGym’s search API defaults to the ClueWeb22-B corpus given the higher alignment with the Researchy Questions benchmark.
4.2 System-level Evaluation
Table 2: Comparison of deep research systems on the Researchy Questions test set using (i) each system’s original commercial search API and (ii) DeepResearchGym’s search API (ours). Scores are judged by gpt-4.1-mini-2025-04-14. Systems marked with * are not tailored for long-report generation.
@@ 4.2 System-level Evaluation
System Commercial Ours Commercial Ours Commercial Ours
@@
The evaluation also reveals a trade-off in commercial systems, which tend to produce strong narrative quality but sometimes with reduced citation precision. Manual inspection shows two recurring issues: (1) citations are often used to support broad sections instead of specific claims, and (2) some cited URLs cannot be fully crawled. This points to a tension between narrative coherence and precise evidence anchoring in current designs. For these citation metrics, note that systems that do not present references receive a Faithfulness scores of 0. Appendix F presents additional runs using different judges and search corpora, and Appendix G presents a qualitative analysis on failure modes.
@@ 4.3 Query-level Analysis
Figure 2: Query-level correlation across metrics, when changing between commercial search and our API.
To further investigate the consistency of system performance across individual queries, we conducted a fine-grained analysis comparing results obtained under each system’s original retrieval API, against those from DeepResearchGym’s API, focusing only on the systems geared towards long-report generation and explicit references. Figure 2 presents scatter plots of per-query scores across our three evaluation axes.
@@ 4.4 Human Evaluation
To validate our automatic evaluation protocol, we conducted a human study over 210 queries with their corresponding reports. For each query, three annotators (drawn from a pool of seven co-authors) compared two system outputs and selected the better one with respect to informativeness, coherence, and factual accuracy. The study was conducted double-blind, with randomization of system assignment and report order, and ties were disallowed to enforce binary preferences.
Inter-annotator reliability was high, with an average pairwise Cohen’s $\kappa$ of 0.87. Agreement between LLM-based automatic judgments and human preferences was similarly strong: $\kappa=0.72$ for KPR, $0.86$ for both citation precision and recall, $0.89$ for clarity, and $0.84$ for insightfulness. The KPC metric was excluded due to insufficient non-tied comparisons. Across dimensions, the same relative system ranking was observed under human and automatic judgments, confirming that our LLM-as-a-judge protocol reflects human preferences.
5 Case Study: DeepResearchGym API for Search Agent Training
Beyond evaluation, another application of DeepResearchGym’s search API is the cost-effective training of agentic search systems. Training search agents through reinforcement learning needs thousands of search API calls. For instance, an experiment with 10,000 queries, 16 trajectories per query, and 4 searches per trajectory results in 640,000 API calls - costing up to US\$640 with SERPER or US\$5,000 with Tavily. For multiple experimental runs, this rapidly becomes expensive, besides being subject to the evolving API behavior.
We empirically show that agents trained using DeepResearchGym’s search API generalize to both unseen benchmarks and commercial search environments, validating the framework as a practical training environment. We focus on short-form question answering rather than long-form report generation, as this setup is more efficient to train and rewards are easier to establish, while maintaining the core search behaviors.
5.1 Agent Training and Inference
We train a search agent with Qwen3-1.7B (DBLP:journals/corr/abs-2505-09388) as the backbone LLM, equipped with three actions: search, summarize context, or answer (DBLP:journals/corr/abs-2510-06534). Training follows GRPO (DBLP:journals/corr/abs-2402-03300), using LLM-as-a-judge soft-match rewards between agent responses and ground-truth answers. We compare two configurations: First, we train with commercial search API (Serper) on synthetic data from AFM-Web-Agent (DBLP:journals/corr/abs-2508-13167). Second, we synthesize training queries by adapting existing approaches (DBLP:journals/corr/abs-2508-07976; DBLP:journals/corr/abs-2506-10055; DBLP:journals/corr/abs-2508-13167) to be grounded over ClueWeb22 rather than the live web, and train with DeepResearchGym’s API. Both configurations train on 5,000 queries.
At inference time, we evaluate under two scenarios. First, to isolate the effect of training infrastructure, both agents are evaluated on standard benchmarks, using Serper to query the live web regardless of which API they were trained with. Second, to provide an in-distribution comparison, we evaluate each agent using its respective training API on a curated subset of standard benchmarks that is covered by ClueWeb22, ensuring both agents have access to relevant information. Further details can be found in Appendix H.
5.2 Results
Table 3: Search agent performance before and after reinforcement learning with different search engines.
Method Search Train Search Infer GAIA WebWalker HLE ClueWeb-Test
Qwen3-1.7B (DBLP:journals/corr/abs-2510-06534) (no training) Serper 8.7 19.1 6.2 16.9
Qwen3-1.7B (DBLP:journals/corr/abs-2510-06534) (no training) ClueWeb - - - 13.2
Qwen3-1.7B + RL (Open Web) Serper Serper 18.4 35.4 6.9 24.5
Qwen3-1.7B + RL (DeepResearchGym) ClueWeb ClueWeb - - - 18.8
Qwen3-1.7B + RL (DeepResearchGym) ClueWeb Serper 20.3 29.7 7.0 22.6
Table 3 presents pass@1 success rates on three benchmarks: GAIA (DBLP:conf/iclr/MialonF0LS24), WebWalkerQA (DBLP:conf/acl/0007Y0WXFZ0ZXH25), and HLE (phan2025humanitysexam). ClueWeb-Test is a smaller 53-query subset of these benchmarks answerable using ClueWeb22, for which we report pass@4. Although this subset is small, both setups exhibit comparable relative improvements, which supports the use of DeepResearchGym as a stable environment for probing training dynamics. Furthermore, the ClueWeb-trained agent generalizes effectively to commercial search at inference, matching or exceeding the Serper-trained baseline on GAIA and HLE, though trailing on WebWalkerQA. While an expected performance gap exists between search backends, these results confirm that DeepResearchGym enables cost-effective and reproducible training, where learned search strategies transfer across both engines and evaluation distributions.
@@ 6 Conclusion and Future Work
DeepResearchGym offers a reproducible sandbox for developing and benchmarking deep research systems, providing a stable cost-effective alternative to commercial search APIs. By anchoring retrieval to high-quality web corpora, our framework enables controlled experimentation across diverse use cases, from agent training to systematic evaluation of report generation systems.
@@
Results demonstrate that DeepResearchGym serves as a reliable research-grade complement to commercial retrieval infrastructures. Evaluation-wise, Systems maintain comparable performance when transitioning from proprietary APIs to our transparent environment, confirming preserved retrieval fidelity. Beyond evaluation, our case study shows that agents trained exclusively within DeepResearchGym generalize effectively to commercial search at inference time, validating the framework as a practical training environment. By isolating system behavior from fluctuating retrieval conditions and API expenses, DeepResearchGym provides a foundation for reproducible, accessible research in deep research systems.
@@
Future extensions to DeepResearchGym can expand the coverage to recent web corpora, such as newer FineWeb crawls, enabling evaluation of time-sensitive queries and emerging topics. Moreover, the integration of domain-specific benchmarks may further support assessment in high-stakes contexts such as healthcare or law, where retrieval precision and factual reliability are critical.
Reproducibility Statement: DeepResearchGym is explicitly designed to enhance reproducibility in deep research systems by providing a free and transparent search API over public web corpora (ClueWeb22 and FineWeb). All code for the framework, including retrieval pipelines, evaluation scripts, and experimental configurations, are available as open-source software. The embedding model used in our dense retrieval pipeline is publicly available, and instructions for replicating experiments are provided. The retrieval API has also been made available as an online service, and Appendix A presents an analysis of requests that have been submitted. We do acknowledge that the automated evaluation used in the manuscript leverages LLMs as judges, which introduces some inherent variability. However, this limitation is standard in current research practice and is documented in the methodology. Also, we used a proprietary LLM to support the evaluation, which brings further problems in terms of reproducibility. Still, this choice was made in support of maximizing the evaluation effectiveness, although future enhancements can perhaps consider open LLMs instead. The URLs for the public search API and the GitHub repository containing the source code will be released online.
LLM Usage Statement: We used LLMs to assist with phrasing, improve clarity and readability, and help to summarize longer sections. However, these tools were only used for language refinement, and did not contribute to the research content or results.
Appendix A Query Log Analysis
Figure 3: Query length distribution (left), and frequency of the 25 most common first-words in the query log (right).
Figure 4: Query frequency distribution (left), and query type distribution (right).
We made publicly available a standardized search API that researchers could integrate into their deep research systems, as an alternative to commercial web search APIs commonly used in prior work. As described in Section 3.1.3, one of the endpoints takes a query, and returns $n$ documents from either FineWeb, ClueWeb-B, or ClueWeb-A, as chosen by the user. To assess the initial adoption of this service over the first four months, we analyzed a query log comprising roughly 12 million queries submitted by 384 unique IP addresses across 13 countries.
We first classified a representative sample of queries using a standard web search taxonomy that considers informational, navigational, and transactional intents. Prior studies of general-purpose search engines report that around 50% of queries are informational (10.1145/792550.792552). In contrast, our logs reveal a markedly higher proportion, with approximately 90% of queries being classified as as informational by gpt-4.1-mini-2025-04-14. This suggests that users primarily employed the service for knowledge-seeking purposes, as expected from deep research systems.
Figure 3 (left) shows the query length distribution. The average query length is 6.7 words, which is considerably higher than typical web search queries, generally reported between 3 and 4 words (10.1145/1507509.1507511; roy2016syntactic). This further suggests that queries are more complex, consistent with the goals of a deep research system. In turn, Figure 3 (right) displays the top-25 most frequent first words in queries, highlighting a dominance of WH-questions (e.g., who, when) and geo-spatial terms (e.g., location, country), another pattern consistent with information-seeking behavior.
Finally, Figure 4 (left) shows the query frequency distribution, which follows an expected Zipfian distribution (XieOHallaron2002). Singleton queries account for roughly 27% of all queries, representing a large fraction of unique interactions. In contrast, the long tail of repeating queries contributes substantially to overall query volume, which may include e.g. benchmarking or systematic evaluation of models. Looking into the queries themselves, Figure 4 (right) further characterizes them by MS-MARCO types (DBLP:conf/nips/NguyenRSGTMD16), with PERSON, DESCRIPTION, and LOCATION queries being the most frequent.
Appendix B Researchy Questions